# Machine-Learning Estimation of Energy Fractions in MHD Turbulence Modes

### v1

@@ Abstract
Magnetohydrodynamic (MHD) turbulence plays a central role in many astrophysical processes in the interstellar medium (ISM), including star formation, heat conduction, and cosmic-ray scattering. MHD turbulence can be decomposed into three fundamental modes—fast, slow, and Alfvén—each contributing differently to the dynamics of the medium. However, characterizing and separating the energy fractions of these modes was challenging due to the limited information available from observations. To address this difficulty, we use 3D isothermal and multiphase MHD turbulence simulations to examine how mode energy fractions vary under different physical conditions. Overall, we find that the Alfvén and slow modes carry comparable kinetic-energy fractions and together dominate the turbulent energy budget in multiphase media, while the fast mode contributes the smallest fraction. Relative to isothermal conditions, multiphase simulations exhibit an enhanced fast-mode energy fraction. We further introduce a machine–learning–based approach that employs a conditional Residual Neural Network to infer these fractions directly from spectroscopic data. The method leverages the fact that the three MHD modes imprint distinct morphological signatures in spectroscopic maps owing to their differing anisotropies and compressibilities. Our model is trained on a suite of isothermal and multiphase simulations covering typical ISM conditions. We further demonstrate that our machine learning model can robustly recover the mode fractions from spectroscopic observables, achieving mean absolute errors of approximately 0.05 for seen data and 0.1 for unseen data.
@@
1 Introduction
Magnetohydrodynamic (MHD) turbulence is a fundamental component of the interstellar medium (ISM), shaping its dynamical, thermal, and chemical evolution (Linde et al., 1998; Goldreich & Sridhar, 1995; Ng et al., 2003; Hu, 2025a). Through the cascade of energy from large to small scales, it governs the structure and dynamics of the multiphase ISM and thereby regulates star formation (Mac Low & Klessen, 2004; McKee & Klein, 2012; Federrath & Klessen, 2012). MHD turbulence mediates the transport of heat, momentum, and mass between different ISM phases (Draine, 2009) and strongly influences the propagation and acceleration of cosmic rays (Jokipii, 1966; Yan & Lazarian, 2002, 2008; Xu & Yan, 2013; Hu et al., 2025). These processes collectively establish MHD turbulence as a central agent in linking microphysical plasma processes to the large-scale evolution of galaxies.
Over the past few decades, theoretical advances have greatly deepened our understanding of MHD turbulence (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999), while numerical simulations have provided crucial tests of these theoretical frameworks (Cho & Vishniac, 2000). In particular, Cho & Lazarian (2003a) analyzed the scaling and anisotropy of compressible MHD turbulence by decomposing it into three fundamental eigenmodes: fast, slow, and Alfvén modes. The Alfvén and slow modes follow a Kolmogorov-like $k^{-5/3}$ spectrum with the scale-dependent Goldreich–Sridhar anisotropy (Goldreich & Sridhar, 1995), whereas fast modes exhibit an isotropic $k^{-3/2}$ spectrum in subsonic condition or $k^{-2}$ spectrum in supersonic condition (Cho & Lazarian, 2003a; Kowal & Lazarian, 2010; Hu et al., 2022; Hou et al., 2025).
Isotropic fast modes play a dominant role in scattering cosmic rays through gyro-resonant interactions, whereas the highly anisotropic Alfvén and slow modes are less efficient due to the averaging effects over many eddies during gyro-resonance (Yan & Lazarian, 2002, 2004). Consequently, the relative energy fraction distribution among the three modes is a key factor in determining cosmic-ray transport in the ISM, yet constraining this distribution observationally remains highly challenging.
@@ 1 Introduction
This paper is organized as follows. § 2 presents the theoretical framework for mode decomposition and explains how morphological features—intensity, centroid velocity, and channel maps—encode information about MHD turbulence modes in both isothermal and multi-phase ISM conditions. § 3 describes the MHD turbulence simulations employed in this study and outlines the ResNet-based neural network architecture and training methodology. § 4 presents the results of our numerical experiments. § 5 compares our findings with previous work and discusses future prospects of the machine-learning approach. Finally, § 6 summarizes the key results and conclusions.
2 Theoretical consideration
2.1 Properties of MHD modes
In ideal compressible MHD turbulence, fluctuations can be decomposed into three distinct modes: the incompressible Alfvén mode and two compressible modes, commonly referred to as the fast and slow modes. Cho & Lazarian (2002, 2003b) introduced a decomposition method to study the anisotropy of three modes. The corresponding equations defining the basis vectors for the mode decomposition are:
$\hat{\xi}_{\rm s}\propto\left(1+\frac{\beta}{2}-\sqrt{D}\right)k_{\bot}\hat{k}_{\bot}+\left(-1+\frac{\beta}{2}-\sqrt{D}\right)k_{\parallel}\hat{k}_{\parallel}$, (1)
$\hat{\xi}_{\rm f}\propto\left(1+\frac{\beta}{2}+\sqrt{D}\right)k_{\bot}\hat{k}_{\bot}+\left(-1+\frac{\beta}{2}+\sqrt{D}\right)k_{\parallel}\hat{k}_{\parallel}$,
$\hat{\xi}_{\rm a}\propto\hat{k}_{\parallel}\times\hat{k}_{\bot}$,
where $D=(1+\beta/2)^{2}-2\beta\cos^{2}\theta$, $\beta=2M_{A}/M_{S}$, $\theta$ is the angle between the wavenumber $\hat{\mathbf{k}}$ and the mean magnetic field $\mathbf{B_{0}}$, and $\boldsymbol{\xi}$ is the displacement vector, with the subscript denoting the corresponding mode: Alfvénic (a), Fast (f), or Slow (s) (see Fig. 1).
@@ 2.1 Properties of MHD modes
Figure 2: Illustration of three MHD turbulence modes after mode decomposition. Panel (a): isothermal simulation min05 ($\beta=0.02$). Panel (b): multi-phase simulation B6v5 ($\beta=0.0061$).
@@ 3 Methodology
3.1 3D isothermal MHD turbulence simulation
@@
Kinetic energy is solenoidally injected at wavenumber 2 to produce a Kolmogorov spectrum. We continuously drive turbulence and dump the data until the turbulence gets fully developed at one sound crossing time. The simulation is grid into $512^{3}$ cells. The simulations are characterized by the sonic Mach number $M_{s}$ and Alfvén Mach number $M_{A}$. In this work, we refer to the simulations in Table 1 by their model name or key parameters.
Table 1: Overview of isothermal simulation runs in this work.
Model $M_{S}$ $M_{A}$ $\beta$
min01 11.05 0.22 0.0008
min02 11.02 0.43 0.003
min03 10.86 0.69 0.008
min04 10.50 0.81 0.010
min05 10.22 1.07 0.020
min06 2.17 0.51 0.110
3.2 3D Multiphase ISM simulations
@@
Similarly, the simulation box size is 100 pc, with turbulence driven solenoidally at a peak wavenumber of 2. The computational domain is discretized on a uniform $512^{3}$ cells. The initial conditions include a uniform number density field of $n=3\text{cm}^{-3}$ and a uniform magnetic field aligned along the y-axis. We choose $B\approx 3~{\rm\mu G}$ and $B\approx 6~{\rm\mu G}$ based on the Zeeman observation (Crutcher, 2012). Two values of velocity dispersion are included: $\sigma_{v}=2.5$ and $\sigma_{v}=5~\text{km s}^{-1}$. These parameters are summarized in Table.2. For details, we refer to Hu (2025a).
Table 2: Overview of Multi-Phase simulation runs in this work.
Model $B[\mu G]$ $\sigma_{v}[km/s]$ $\beta$
B6v2.5 6.0 2.5 0.067
B3v2.5 3.0 2.5 0.175
B6v5 6.0 5.0 0.061
3.3 Synthetic spectroscopic observation
To create synthetic spectroscopic observations as our ML training input, we utilize velocity and density fields derived from MHD simulations to construct PPV cubes. Following early work on $M_{A}$ prediction (Schmaltz et al., 2025), we generate intensity maps, velocity centroids, and velocity channel maps as our model training morphological features.
@@ 3.3 Synthetic spectroscopic observation
$C(x,y)=\frac{\int\rho(x,y,v_{\rm los})v_{\rm los}dv_{\rm los}}{\int\rho(x,y,v_{\rm los})dv_{\rm los}}$ (6)
Velocity channel maps, $p(x,y)$, are derived by integrating the PPV cube over a narrowly defined velocity range. These channel maps are predominantly affected by velocity fluctuations due to the phenomenon known as velocity caustics (Lazarian & Pogosyan, 2000; Hu & Lazarian, 2023). The expression for a thin velocity channel is given by:
@@
These three maps contain complementary information on the density and velocity fields. The intensity map primarily traces the integrated density along the line of sight, and therefore carries information mainly about the underlying density structure. In contrast, the velocity centroid map incorporates both density and velocity information. It provides an intensity-weighted measure of the line-of-sight velocity; thus, it is sensitive to the velocity field when the density fluctuations are moderate. The velocity channel maps, on the other hand, are dominated by velocity fluctuations due to the velocity caustics effects (Lazarian & Pogosyan, 2000; Hu & Lazarian, 2023). In general, the three selected feature maps characterize different proportions of the density and velocity field information.
3.4 Neural Net Architecture
@@ 3.5 Training Strategy
Figure 6: Histograms of the uncertainty $\sigma$ in prediction. $\sigma$ is defined as the absolute difference between the predicted value and the ground truth. Here we demonstrate simulation min01 in the isothermal case and simulation B3v2.5 in the multiphase case. Panel (a)-(b): $\sigma$ distribution when data are included in the training inputs. Panel (c)-(d): $\sigma$ distributions when data are excluded in the training inputs. In each panel, from left to right, represent three modes of MHD turbulence. Blue, green, and red refer to different training inputs.
4 Results
4.1 Mode Decomposition in Multi-phase ISM Simulations
Fig. 3 shows the kinetic‐energy fractions of the three MHD modes as functions of plasma $\beta$ in the multi-phase ISM simulations, compared with those in isothermal simulations. The energy fraction of each mode is defined as the ratio between its kinetic energy and the sum of kinetic energies over all MHD modes, i.e. $f_{i}=E_{i}/(E_{A}+E_{f}+E_{s})$, where $E_{i}=0.5v_{i}^{2}$ and $v_{i}$ is the velocity of the corresponding mode. We find that the Alfvén mode consistently dominates, carrying approximately 40–65% of the total kinetic energy as $\beta$ increases from $10^{-3}$ to $10^{-1}$. The fast mode contributes the smallest fraction, rising from about 5-25% over the same $\beta$ range. The slow-mode fraction remains intermediate between the two. No significant difference is observed between the isothermal and multi-phase cases.
@@ 4.1 Mode Decomposition in Multi-phase ISM Simulations
4.2 Comparison of Model Prediction and Ground Truth
Fig. 5 presents a representative case from the simulation with $B\sim 3~\mu{\rm G}$ and $\sigma_{v}\sim 2.5~{\rm km~s^{-1}}$, illustrating how distinct local structures in the feature maps arise from different fractions of the three MHD modes. The channel, intensity, and centroid maps each display characteristic morphological signatures. In regions where the fast mode dominates—specifically the left–middle and right–middle portions of the map—the centroid field exhibits two prominent, high-amplitude isotropic structures, whereas the corresponding channel map shows two low-intensity features. These features reflect the intrinsically isotropic and compressive nature of the fast mode.
We also compare the ResNet prediction with the ground-truth mode energy fractions in Fig. 5. The model accurately recovers the spatial distribution and relative fraction of the three modes. Moreover, the choice of input—channel, intensity, or centroid map—has only a marginal impact on the reconstruction accuracy.
Next, we evaluate the prediction performance of the proposed conditional ResNet model in both isothermal and multiphase simulations. Fig. 6 shows the histograms of the prediction error, $\sigma$, defined as the absolute difference between the prediction and the ground truth. We consider two evaluation settings: (1) training on all available data and testing on the same (seen) patches, and (2) withholding one patch during training and testing solely on that unseen patch.
For seen data, $\sigma$ remains below 0.02 for all three modes in both isothermal and multiphase simulations, with the fast-mode prediction achieving the highest accuracy (Figs. 6(a) and (b)). For unseen data, $\sigma$ stays within 0.5 in both cases, though the distributions develop broader tails. The mean error in the multiphase case ($\sigma\sim 0.10$–0.16) is higher than in the isothermal case ($\sigma\sim 0.04$–0.09), indicating that the model performs better when predicting mode fractions in isothermal turbulence. Fast-mode predictions consistently outperform those of the Alfvén and slow modes across all scenarios (Figs. 6(c) and (d)). Overall, the error distributions peak near zero in all configurations, demonstrating good consistency and predictive robustness. Among the three modes, the fast mode exhibits the smallest error throughout.
5 Discussion
5.1 Comparison with earlier work

### v2

@@ Abstract
Magnetohydrodynamic (MHD) turbulence plays a central role in many astrophysical processes in the interstellar medium (ISM), including star formation and cosmic-ray transport and acceleration. MHD turbulence can be decomposed into three fundamental modes—fast, slow, and Alfvén—each contributing differently to the dynamics of the medium. However, characterizing and separating the energy fractions of these modes was challenging due to the limited 2D information available from observations. To address this difficulty, we use 3D isothermal and multiphase MHD turbulence simulations to examine how mode energy fractions vary under different physical conditions. Overall, we find that the Alfvén and slow modes carry comparable kinetic-energy fractions and together dominate the turbulent energy budget in multiphase media, while the fast mode contributes the smallest fraction. Relative to isothermal conditions, multiphase simulations exhibit an enhanced fast-mode energy fraction. We further introduce a machine–learning–based approach that employs a conditional Residual Neural Network to infer these fractions directly from spectroscopic data. The method leverages the fact that the three MHD modes imprint distinct morphological signatures in spectroscopic maps owing to their differing contributions to density and velocity fluctuations. Our model is trained on a suite of isothermal and multiphase simulations covering typical ISM conditions. We demonstrate that our machine learning model can recover the mode fractions from spectroscopic observables, achieving mean relative normalized errors of approximately 0 and standard deviation of 0.01 - 0.02 for seen data and 0.1 - 1.8 for unseen data.
@@
I Introduction
Magnetohydrodynamic (MHD) turbulence is a fundamental component of the interstellar medium (ISM), shaping its dynamical, thermal, and chemical evolution (Linde et al., 1998; Goldreich & Sridhar, 1995; Ng et al., 2003; Burkhart et al., 2010; Federrath et al., 2016; Beattie et al., 2020; Ha et al., 2021; Tress et al., 2024; Hu, 2025a; Hu et al., 2025a; Narayan et al., 2025). Through the cascade of energy from large to small scales, it governs the structure and dynamics of the multiphase ISM and thereby regulates star formation (Mac Low & Klessen, 2004; McKee & Klein, 2012; Federrath & Klessen, 2012; Hu & Lazarian, 2023a). MHD turbulence mediates the transport of heat, momentum, and mass between different ISM phases (Draine, 2009; Hu, 2025a) and strongly influences the propagation and acceleration of cosmic rays (Jokipii, 1966; Yan & Lazarian, 2002, 2008; Xu & Yan, 2013; Kempski & Quataert, 2022; Lazarian et al., 2023; Hu et al., 2025b; Hopkins, 2025; Ewart et al., 2026). These processes collectively establish MHD turbulence as a central agent in linking microphysical plasma processes to the large-scale evolution of galaxies.
Over the past few decades, theoretical advances have greatly deepened our understanding of MHD turbulence (Goldreich & Sridhar, 1995; Lazarian & Vishniac, 1999), while numerical simulations have provided crucial tests of these theoretical frameworks (Cho & Vishniac, 2000). In particular, Cho & Lazarian (2003a) analyzed the scaling and anisotropy of compressible MHD turbulence by decomposing it into three fundamental eigenmodes: fast, slow, and Alfvén modes. The Alfvén and slow modes follow a Kolmogorov-like $k^{-5/3}$ spectrum with the scale-dependent anisotropy (Lazarian & Vishniac, 1999), whereas fast modes exhibit an isotropic $k^{-3/2}$ spectrum in subsonic condition or $k^{-2}$ spectrum in supersonic condition (Cho & Lazarian, 2003a; Kowal & Lazarian, 2010; Hu et al., 2022; Hou et al., 2025).
Isotropic fast modes play a dominant role in scattering cosmic rays through gyro-resonant interactions, whereas the highly anisotropic Alfvén and slow modes are less efficient due to the averaging effects over many eddies during gyro-resonance (Yan & Lazarian, 2002, 2004; Hu et al., 2022; Lazarian et al., 2023). Consequently, the relative energy fraction distribution among the three modes is a key factor in determining cosmic-ray transport in the ISM, yet constraining this distribution observationally remains highly challenging.
@@ I Introduction
This paper is organized as follows. § II presents the theoretical framework for mode decomposition and explains how morphological features—intensity, centroid velocity, and channel maps—encode information about MHD turbulence modes in both isothermal and multi-phase ISM conditions. § III describes the MHD turbulence simulations employed in this study and outlines the ResNet-based neural network architecture and training methodology. § IV presents the results of our numerical experiments. § V compares our findings with previous work and discusses future prospects of the machine-learning approach. Finally, § VI summarizes the key results and conclusions.
II Theoretical consideration
II.1 Properties of MHD modes
In ideal compressible MHD turbulence, fluctuations can be decomposed into three distinct modes: the incompressible Alfvén mode and two compressible modes, commonly referred to as the fast and slow modes. Cho & Lazarian (2002, 2003b) introduced a decomposition method to study these three modes. The corresponding basis vectors $\hat{\boldsymbol{\xi}}$ defining the mode decomposition are given by:
$\hat{\boldsymbol{\xi}}_{\rm s}$ $\propto\left(1+\frac{\beta^{\prime}}{2}-\sqrt{D}\right)k_{\bot}\hat{k}_{\bot}+\left(-1+\frac{\beta^{\prime}}{2}-\sqrt{D}\right)k_{\parallel}\hat{k}_{\parallel}$, (1)
$\hat{\boldsymbol{\xi}}_{\rm f}$ $\propto\left(1+\frac{\beta^{\prime}}{2}+\sqrt{D}\right)k_{\bot}\hat{k}_{\bot}+\left(-1+\frac{\beta^{\prime}}{2}+\sqrt{D}\right)k_{\parallel}\hat{k}_{\parallel}$,
$\hat{\boldsymbol{\xi}}_{\rm a}$ $\propto\hat{k}_{\parallel}\times\hat{k}_{\bot}$,
where $D=(1+\beta^{\prime}/2)^{2}-2\beta^{\prime}\cos^{2}\theta$, $\beta^{\prime}=2M_{A}/M_{s}$, and $\theta$ is the angle between the wavevector $\hat{\mathbf{k}}$ and the mean magnetic field $\mathbf{B}_{0}$. Here the subscripts $a$, $f$, and $s$ refer to the Alfvén, fast, and slow modes, respectively (see Fig. 1). $M_{A}$ and $M_{s}$ are the Alfvén and sonic Mach numbers, respectively.
@@ II.1 Properties of MHD modes
Figure 2: Illustration of three MHD turbulence modes’ kinetic energy fraction after mode decomposition. Panel (a): isothermal simulation min05 ($\beta=0.02$). Panel (b): multi-phase simulation B6v5 ($\beta=0.0061$).
@@ III Methodology
III.1 3D isothermal MHD turbulence simulation
@@
Table 1: Overview of isothermal simulation runs in this work. $\beta=2(M_{A}/M_{s})^{2}$ is the plasma beta. $\delta\rho/\langle\rho\rangle=\sqrt{\langle(\rho-\langle\rho\rangle)^{2}\rangle}/\langle\rho\rangle$ is the density RMS fluctuation normalized to the uniform background density.
Model $M_{s}$ $M_{A}$ $\beta$ $\delta\rho/\langle\rho\rangle$
min01 11.05 0.22 0.0008 0.400
min02 11.02 0.43 0.003 0.521
min03 10.86 0.69 0.008 0.534
min04 10.50 0.81 0.010 0.568
min05 10.22 1.07 0.020 0.573
min06 2.17 0.51 0.110 0.273
Kinetic energy is solenoidally injected at wavenumber 2 and turbulence is naturally evolved to produce a Kolmogorov spectrum. We continuously drive turbulence and dump the data until the turbulence gets fully developed at one sound crossing time. The simulation is grid into $512^{3}$ cells. The simulations are characterized by the sonic Mach number $M_{s}$ and Alfvén Mach number $M_{A}$. In this work, we refer to the simulations in Table 1 by their model name or key parameters.
III.2 3D Multiphase ISM simulations
@@
[…]
Table 2: Overview of multi-phase simulation runs in this work. Here $\beta$ is calculated from the mean $M_{A}$ amd $M_{s}$.
[…]
B6v2.5 6.0 2.5 0.067 0.379
B3v2.5 3.0 2.5 0.175 0.551
B6v5 6.0 5.0 0.061 0.451
III.3 Synthetic spectroscopic observation
To create synthetic spectroscopic observations as our ML training input, we utilize velocity and density fields derived from MHD simulations to construct PPV cubes. Following early work on $M_{s}$ prediction (Schmaltz et al., 2025), we generate intensity maps, velocity centroids, and velocity channel maps as our model training morphological features.
@@ III.3 Synthetic spectroscopic observation
$C(x,y)=\frac{\int\rho(x,y,v_{\rm los})v_{\rm los}dv_{\rm los}}{\int\rho(x,y,v_{\rm los})dv_{\rm los}}.$ (6)
Velocity channel maps, $p(x,y)$, are derived by integrating the PPV cube over a narrowly defined velocity range. These channel maps are predominantly affected by velocity fluctuations due to the phenomenon known as velocity caustics (Lazarian & Pogosyan, 2000; Hu & Lazarian, 2023b). The expression for a thin velocity channel is given by:
@@
These three maps contain complementary information on the density and velocity fields. The intensity map primarily traces the integrated density along the line of sight, and therefore carries information mainly about the underlying density structure. In contrast, the velocity centroid map incorporates both density and velocity information. It provides an intensity-weighted measure of the line-of-sight velocity; thus, it is sensitive to the velocity field when the density fluctuations are moderate. The velocity channel maps, on the other hand, are dominated by velocity fluctuations due to the velocity caustics effects (Lazarian & Pogosyan, 2000; Hu & Lazarian, 2023b). In general, the three selected feature maps characterize different proportions of the density and velocity field information.
III.4 Neural Net Architecture
@@ III.5 Training Strategy
Figure 6: Histograms of the relative prediction error $\sigma$, which is defined as the difference between the predicted ($f^{\rm pred}$) and true $f^{\rm ture}$ energy fractions, normalized by the ground-truth value for each mode, i.e., $\sigma=|f_{i}^{\rm pred}-f_{i}^{\rm true}|/f_{i}^{\rm true}$, where the subscript $i=A,S,{\rm or}~F$ denotes the Alfvén, slow, and fast modes, respectively. Here we demonstrate simulation min01 in the isothermal case and simulation B3v2.5 in the multiphase case. Panel (a)-(b): $\sigma$ distribution when data are included in the training inputs. Panel (c)-(d): $\sigma$ distributions when data are excluded in the training inputs. Each panel, from left to right, represents Alfvén, slow, and fast modes, respectively. Blue, green, and red refer to different training inputs. ”std” means the standard deviation.
IV Results
IV.1 Mode Decomposition in Multi-phase ISM Simulations
Fig. 3 shows the kinetic‐energy fractions of the three MHD modes as functions of plasma $\beta$ in the multi-phase ISM simulations, compared with those in isothermal simulations. The energy fraction of each mode is defined as the ratio between its kinetic energy and the sum of kinetic energies over all MHD modes, i.e. $f_{i}=E_{i}/(E_{A}+E_{f}+E_{s})$, where $E_{i}=0.5v_{i}^{2}$ and $v_{i}$ is the velocity of the corresponding mode. We find that the Alfvén mode consistently dominates, carrying approximately 40–65% of the total kinetic energy as $\beta$ increases from $10^{-3}$ to $10^{-1}$. The fast mode contributes the smallest fraction, rising from about 5-25% over the same $\beta$ range. The slow-mode fraction remains intermediate between the two. Relative to isothermal conditions, multiphase simulations exhibit an enhanced fast-mode energy fraction of about 20 - 30%.
@@ IV.1 Mode Decomposition in Multi-phase ISM Simulations
IV.2 Comparison of Model Prediction and Ground Truth
[…]
We also compare the ResNet prediction with the ground-truth mode energy fractions in Fig. 5. The model accurately recovers the spatial distribution and relative fraction of the three modes. Moreover, the choice of input—channel, intensity, or centroid map has only a marginal impact on the reconstruction accuracy.
Next, we evaluate the prediction performance of the proposed conditional ResNet model in both isothermal and multiphase simulations. Fig. 6 shows the histograms of the relative prediction error, $\sigma$, defined as the difference between the predicted ($f^{\rm pred}$) and true ($f^{\rm ture}$) energy fractions, normalized by the ground-truth value for each mode, i.e., $\sigma=|f_{i}^{\rm pred}-f_{i}^{\rm true}|/f_{i}^{\rm true}$, where the subscript $i=A,S,{\rm or}~F$ denotes the Alfvén, slow, and fast modes, respectively. We consider two evaluation settings: (1) training on all available data and testing on the same (seen) patches, and (2) withholding one patch during training and testing solely on that unseen patch.
For seen data, $\sigma$ remains below 0.3 for all three modes in both isothermal and multiphase simulations, with the fast-mode prediction achieving the highest accuracy (Figs. 6(a) and (b)). For unseen data, $\sigma$ generally stays within 2 in both cases, though the distributions develop broader tails. The mean error in the isothermal case ($\sigma\sim 0.10–1.79$) is higher than in the multiphase case ($\sigma\sim 0.35–0.83$), indicating that the model performs better when predicting mode fractions in multiphase turbulence. Alfvén-mode predictions consistently outperform those of the fast and slow modes across all scenarios (Figs. 6(c) and (d)). Overall, the relative error distributions generally peak at zero across all configurations, except for fast-mode predictions on unseen data, indicating good consistency and predictive robustness of the model. Among the three modes, the Alfvén mode exhibits the smallest errors, whereas the fast mode remains the most unstable. This behavior may be partially attributed to the isotropic nature of the fast mode, which leads to less distinct morphological structures in the sliced feature maps.
V Discussion
V.1 Comparison with earlier work